OpenStudy (hpfan101):

\[\lim_{x \rightarrow \infty}=\tan^{-1} (e^x)\]

OpenStudy (anonymous):

same as \[\lim_{x \rightarrow \infty}=\tan^{-1} (x)\]

OpenStudy (hpfan101):

Ok, thank you. You just did natural log to get rid of "e" right?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

as \(x\to \infty\) you have \(e^x\to\infty\)

OpenStudy (anonymous):

and arctangent is continuous, so if you want to be formal you can take the limit inside, but really it is just that the input is going to infinity

OpenStudy (hpfan101):

Oh, I see. Thanks!